# Integrating Graceful Degradation and Recovery through Requirement-driven Adaptation

### 18 Jan 2024 v1

@@ V-C Runtime Adaptation as a MILP Formulation
Problem V.1
Graceful Degradation as MILP: Given the current requirement $\varphi(\nu(\textbf{p}))$ and its weaker version to be found, $\varphi(\nu^{\prime}(\textbf{p}))$, transition system $T=(\mathcal{Q},\mathcal{A},\delta,\mathcal{Q}_{i})$, and state sequence $q^{t}=q_{0},\dots,q_{t}$ representing the signal observed from the environment so far, compute:
@@
$s.t.~{}$ $~{}\textbf{s}{}_{i}=q_{i}\ \text{for }i\leq t$ (17)
$~{}\textbf{s}{}_{i}=\delta(s_{i-1},\mathbf{a}_{i-1})\ \text{for }t<i\leq t+N$ (18)
$~{}(0,\nu^{\prime}(\textbf{p}))\in d(\textbf{s}{},\varphi)$ (19)
Note that in Eq. (17), the past signal is defined up until the current time. Eq. (18) involves the prediction of future signals from the next time step to time $t+N$ using the transition system $T$ that encodes an environment model. $N\in\mathbb{N}$ is a finite predictive horizon inferred from the STL requirement $\varphi$. Eq. (19) guarantees that the new requirement is within the defined range of requirements by restricting the parameters of the PSTL formula to the validity domain.
@@
Problem V.2
Recovery MILP Formulation: Given the current requirement $\varphi(\nu(\textbf{p}))$ and it strengthened version to be found, $\varphi(\nu^{\prime}(\textbf{p}))$, transition system $T=(\mathcal{Q},\mathcal{A},\delta,\mathcal{Q}_{i})$, and state sequence $q^{t}=q_{0},\dots,q_{t}$ representing the signal observed from the environment so far, compute:
@@
$s.t.~{}$ $~{}\textbf{s}{}_{i}=q_{i}\ \text{for }i\leq t$ (21)
$~{}\textbf{s}{}_{i}=\delta(s_{i-1},\mathbf{a}_{i-1})\ \text{for }t<i\leq t+N$ (22)
$~{}(0,\nu^{\prime}(\textbf{p}))\in d(\textbf{s}{},\varphi)$ (23)
@@
VI Implementation
Figure 3: A screenshot of the UUV simulator.
VI-A Simulator
To illustrate our proposed requirement adaptation approach, we have developed a prototype implementation. Note that I will publish all the artifacts, including controller code, models, and experimental data once the paper is accepted. The implementation is based on SUAVE [3], an unmanned underwater vehicle that supports the customization of self-adaptation logic. It uses a ROS2-based platform and has implemented a pre-defined mission that detects, follows, and inspects a pipeline on the seabed.
To evaluate our proposed approach, we randomly injected abnormal environmental events throughout the operation of the UUVs. We listed two failures that we investigated in our case studies along with their degradation and recovery actions below:
@@ VI-A Simulator
  • Loss of Visual Contact with Underwater Pipeline: The low water visibility makes it difficult for the UUV to detect and follow the underwater pipeline. When it occurs during the undersea pipeline inspection, it may result in the UUV losing sight of the pipeline and requiring it to dive deeper in the short term, regain the visual line of sight, and continue the inspection progress. The change in the water visibility is injected during the simulation randomly.
@@
  • Thruster Failure: Thruster failure causes the engine of the UUV to provide partial propulsion, move erratically, or shut down completely. Similarly, thruster failures are modeled and injected in a stochastic manner.
VI-B Environment Model
@@
Our framework leverages a model of the environment during the adaptation process. For the UUV systems, the environmental model captures the (1) physical dynamics of the system (2) thruster estimation from the engine, and (3) the estimated coordinates of the seabed pipeline. The dynamics model is used to estimate the velocity of the UUVs using the accelerating or decelerating actions. The thruster estimation is based on the configuration of the thrusters (i.e. which one is turned on or off). Lastly, the estimated coordinates of the seabed pipeline are based on information received via onboard sensors. The coordinates information is also used to keep track of the inspection progress and guide the diving operation toward the pipeline.
@@
All aspects of the environmental model are specified in the MiniZinc modeling language, translated to linear constraints in MILP, and solved during the runtime adaptation process.
VII Evaluation
@@
This section describes the evaluation of our proposed approach. We presented the following:
@@
  • RQ1: Does our requirements-driven adaptation approach achieve a higher overall system utility than existing state-of-the-art approaches?
@@
To investigate these RQs, we implemented our requirement adaptation framework for a UUV on an underwater pipeline inspection mission based on the SUAVE exemplar [3].
VII-A Experimental Design
We conducted a set of experiments to evaluate the proposed adaptation approach through comparison against with state-of-the-art self-adaptation method in TOMASys [2].
In the case that the MILP solver is unable to return a feasible solution, it returns UNSAT, indicating unsatisfiability. It may happen under certain environmental disruptions where it is impossible to identify an STL requirement that is stronger than the minimal requirement, in which case the planner maintains the current configuration of the system without adaptation.
To ensure that our adaptation approach generalizes across various scenarios, we randomly generated different system setups and failure scenarios, including the number of sensors, throughput, and energy configurations of each sensor, duration of the simulation, location of the underwater pipeline, and the starting location of the vehicle; in addition, we also varied how much the throughput degrades and the duration of the failure, the duration of thruster failures, and how much the water visibility fluctuates.
@@ VII-A Experimental Design
To compare our proposed method and the baseline approach, we use the cumulative utility as the metric, which is measured by calculating the robustness of the minimal requirement with respect to the signal trace during the recovery process. The reason why we chose to measure the satisfaction of the minimal requirement is that it is the lower bound of the instantiated STL formula under which the system feature ceases to function and provides useful utility. We assume that the robustness metric is a good medium to reflect the desirability of the system behavior. The robustness is measured upon the start of the degradation events and ends upon the satisfaction of the optimal requirement ($\varphi_{opt}$). In the case that the optimal requirement is never reached due to persistent environmental disruptions, the cumulative robustness measurement would continue until the end of the simulation life cycle.
@@
Before we conducted the experiments for the case studies, we proposed the following hypothesis to be tested:
@@
  • H1: Our adaptation approach results in a higher cumulative utility throughout degradation and recovery processes than the existing method.
@@
  • H2: Our adaptation approach results in a higher run-time overhead but it does not disrupt normal system operations.
@@
All our experiments were run on a Ubuntu desktop machine with 16 GB RAM and 6-core Intel Core i5, along with NVIDIA GeForce RTX 3060 graphics card.
VII-B Seabed Pipe Inspection Case Study
Recall the example mentioned in Section II, where an Unmanned Underwater Vehicle (UUV) is conducting a mission to inspect underwater pipes on the seabed. The mission is subject to environmental uncertainties like low water visibility events, in which case the vehicle dives deeper towards the pipeline, and thruster failures, in which case the vehicle reduces speed or floats back up to the surface of the water.
The optimal STL formula $\varphi_{opt}$ and the minimal requirement $\varphi_{min}$ are specified as follows to capture the diving requirement due to water visibility and speed requirement that helps the vehicles complete the mission with a reasonable rate:
@@ VII-B Seabed Pipe Inspection Case Study
$\begin{split}&\varphi_{visibility\_opt}:\Box_{[0,1]}(\texttt{visibility}<20\Rightarrow\\ &\quad\quad\quad\quad\quad\Diamond_{[0,5]}\texttt{distance\_to\_pipeline}<10)\end{split}$ (24)
@@
$\begin{split}&\varphi_{visibility\_min}:\Box_{[0,1]}(\texttt{visibility}<20\Rightarrow\\ &\quad\quad\quad\quad\quad\Diamond_{[0,15]}\texttt{distance\_to\_pipeline}<10)\end{split}$ (25)
$\begin{split}&\varphi_{thruster\_opt}:\Box_{[0,1]}(\texttt{thrust}>100)\end{split}$ (26)
$\begin{split}&\varphi_{thruster\_min}:\Box_{[0,1]}(\texttt{thrust}>50)\end{split}$ (27)
We randomly generated 100 randomized scenarios based on Section VII-A. We run all the scenarios using both the baseline and proposed requirement-driven adaptation approaches.
@@
Figure 4: Cumulative robustness for visibility and thruster monitor features for the pipeline inspection case study
VII-C Experiment Result Discussions
@@
In Fig 4, charts (A) and (B) showed the cumulative robustness values for both the visibility and thruster monitor feature while the system is encountering failures including thruster failure and low water visibility events. The results show that our requirement adaptation achieves a higher cumulative robustness value than the baseline self-adaptive approach, in the case of the water visibility monitor, it achieves an approximately 2-fold increase in the cumulative robustness (with a 26.9 increase in robustness). In the case of the thruster monitor feature, our approach has a 24.3 lead in cumulative robustness.
@@
The SUAVE artifact [3] also comes with a set of metrics that measures the quality of the mission completion, we reused some of the metrics, namely, the distance of the pipeline inspected, and observed that our approach, on average, can inspect 86.10 meters of the pipeline while the baseline approach can only inspect 49.63 meters, an increase of about 74%. However, the standard deviation of our approach (36.68) is slightly worse than the baseline approach (23.30) meaning that the actual performance fluctuates more across various failure scenarios in the 100 experiment runs.
@@
There are two main reasons for the increase in robustness and quality of the mission completion. First, it’s the dynamic nature of the action planning, for example, TOMASys relies on a design time fixed actions to address each scenario. To satisfy the objective of regaining contact with the pipeline after losing visuals, TOMASys simply has a predefined action to dive several meters down the seabed, whereas our approach leverages the environmental model pruning a space of actions to search for the optimal strategy. Second, it’s the ability to flexibly adapt goals in different situations, namely, adjust the goals when they aren’t attainable or have potential for improvements.
VII-D Performance Overhead
@@
Since the requirement adaptation approach requires the use of MILP to search for requirements and plan actions at runtime, it has a higher runtime overhead than our baseline approach which uses actions that are defined at design time. We measured the average overhead across both features as 0.35 seconds. We have not observed noticeable delays or disruptions to our UUV during the operation, as the controller for the ArduSub was running at 2Hz, resulting in a window of 0.5 seconds for each cycle of the controller update.
@@
VII-E Threats to Validity
Our experiments are subject to threats of both internal validity and external validity.
Internal Validity: The sampled configuration space is partial and may not capture some exceptional scenarios. However, we have mitigated selection biases by randomizing the configuration parameter, for example, at which point failures occur, the initial location of the UUV, etc.
@@ VII-E Threats to Validity
External Validity: The overhead of the simulation is application-specific and hardware-dependent. More restrictive hardware or outdated MILP solver may increase the runtime overhead of our framework.
VIII Related Work
@@
Graceful Degradation and Recovery. Graceful degradation is the ability of a system to maintain some acceptable level of functionality even when a significant portion of the system is rendered inoperable due to environmental disturbances. Recovery refers to the ability of the system to restore to a safe or desired state after a failure. Both areas are well-studied in control systems and software applications. A number of degradation frameworks have been developed for different applications and domains, such as adaptive cruise control [9], autonomous drones [10] software user interfaces [11], industrial control systems [12], and design patterns for degradation[13]. There is also a large body of work on system recovery. [14] proposed an STL-based resilient framework that enables analysis of trade-offs between time-to-recovery and durability in a cruise control system using multi-objective optimization, while [15] and [16] propose checkpointing techniques (i.e. storing system traces with a finite horizon leading up to the current state). Lastly, [17] and [18] use system restart and reset to recover the system from a failure. As far as we know, no prior work provides a coordination mechanism to facilitate these two processes on a single system.
@@
Self-adaptive Systems. Self-adaptive systems aim to design systems that are capable of adjusting to changes in the environment. The most influential reference control model in autonomic and self-adaptive systems is the MAPE-K [19] architecture, which stands for Monitor-Analyze-Plan-Execute over a shared Knowledge, on which our resolution architecture is based. Self-healing systems [20] have been proposed by Shaw, et al. to address the problem of coping with fluctuations and uncertainties in the environment. The idea is that the system always keeps a background maintenance process running regardless of the current state of the system and the environment (i.e. garbage collection, optimized network routing, etc). This way, the system will eventually reach a homeostasis state. Our approach draws inspiration from this idea, in that we attempt to ensure that the system aims for the most desirable goal with respect to the current requirement. However, our approach differs from [20] in that we leverage formal requirements specifications in STL, while [20] argues that homeostatic approaches eliminate the need for formal specifications. In addition, the concept of meta-control has been used for design-time analysis. For example, OpenODD [21] [22] has been used as a scenario-based analysis framework for autonomous systems, especially in self-driving cars. TOMASys [2] provides a more comprehensive set of reconfiguration strategies ranging from architectural reconfiguration, to monitor construction, and then to scenario-based definition of adaptation tactics.
Requirement Adaptation. Requirement adaptation has been investigated in the context of self-adaptive systems. The idea of weakening a requirement has been investigated by [10], [12], and [23] to resolve conflicts resulting from feature interactions, violation of environmental assumptions, and safety properties, respectively. RELAX [24] is a temporal logic designed to support expressing requirements that explicitly capture uncertainties about runtime system behavior. RELAX can be used to support self-adaptation in response to environmental changes. However, it has not been explicitly applied to the context of graceful degradation and recovery.
@@
In addition, both [25] and [26] study goal adaptation in the context of changing stakeholder goals or user preferences. However, their proposed methods are designed to be dependent on human inputs, not for automatically adapting to changing environments. [27] focuses on requirement adaptation based on resource constraints, and proposes a 2-tier adaptation framework that (1) first tries to fulfill goals by finding resources or their replacements and (2) if needed, deviates the goal slightly from the original one to compensate for limited resources.
IX Limitation and Future Work

### 08 Apr 2024 v2

@@ 5.3. Runtime Adaptation as a MILP Formulation
Problem 5.1 ().
Graceful Degradation as MILP. Given the current requirement $\varphi(\nu(\textbf{p}))$ and its weaker version to be found, $\varphi(\nu^{\prime}(\textbf{p}))$, transition system $T=(\mathcal{Q},\mathcal{A},\delta,\mathcal{Q}_{i})$, and state sequence $q^{t}=q_{0},\dots,q_{t}$ representing the signal observed from the environment so far, compute:
@@
(13) $s.t.~{}$ $~{}\textbf{s}{}_{i}=q_{i}\ \text{for }i\leq t~{}\land$
(14) $~{}\textbf{s}{}_{i}=\delta(\textbf{s}_{i-1},\mathbf{a}_{i-1})\ \text{for }t<i\leq t+N~{}\land$
(15) $~{}(0,\nu^{\prime}(\textbf{p}))\in d(\textbf{s}{},\varphi)$
There are two decision variables to be solved in this MILP problem: parameters $\nu^{\prime}(\textbf{p})$ and control action sequence a. Note that in Eq. (13), the past signal is defined up until the current time. Eq. (14) involves the prediction of a future signal from the next time step to time $t+N$ using the transition system $T$ that encodes an environment model. $N\in\mathbb{N}$ is a finite predictive horizon inferred from the STL requirement $\varphi$. Eq. (15) guarantees that the new requirement is within the defined range of requirements by restricting the parameters of the PSTL formula to the validity domain.
@@
Problem 5.2 ().
Recovery MILP Formulation. Given the current requirement $\varphi(\nu(\textbf{p}))$ and it strengthened version to be found, $\varphi(\nu^{\prime}(\textbf{p}))$, transition system $T=(\mathcal{Q},\mathcal{A},\delta,\mathcal{Q}_{i})$, and state sequence $q^{t}=q_{0},\dots,q_{t}$ representing the signal observed from the environment so far, compute:
@@
(17) $s.t.~{}$ $~{}\textbf{s}{}_{i}=q_{i}\ \text{for }i\leq t~{}\land$
(18) $~{}\textbf{s}{}_{i}=\delta(\textbf{s}_{i-1},\mathbf{a}_{i-1})\ \text{for }t<i\leq t+N~{}\land$
(19) $~{}(0,\nu^{\prime}(\textbf{p}))\in d(\textbf{s}{},\varphi)$
@@
6. Implementation
6.1. Simulator
To illustrate our proposed requirement adaptation approach, we have developed a prototype implementation [^1] All the code, models, and experimental data are available at https://github.com/sychoo/cps-degradation-recovery based on SUAVE (Silva et al., 2023), an unmanned underwater vehicle that supports the customization of self-adaptation logic. It uses a ROS2-based platform and implements a pre-defined mission that detects, follows, and inspects a pipeline on the seabed. The backend of the simulator uses ArduSub (Zhao et al., 2020), which provides various controller APIs to control the trajectory of the underwater vehicle. The vehicle trajectory and environmental setup can be visualized through the Gazebo simulator (Mengacci et al., 2021) with the UI shown in Fig. 3 where the seabed pipe is indicated with the yellow cylinder, while the UUV is illustrated by the rectangular box to the left side of the pipe. Various environmental entities (i.e., lighting, terrain) are listed on the panel to the right, and can be adjusted as needed.
Features. For evaluation, we implemented two mission-related features in the UUV. When these features are activated, they generate propulsion actions (in the form of velocity vectors) or system reconfiguration actions that override those from the path planner. The implemented features are as follows:
  • Visibility Monitor ensures that the UUV maintains sufficient visibility to the pipelines. To achieve this, it generates actions to enforce the UUV to close in on the pipelines when the visibility is below a safety threshold.
  • Thrust Monitor ensures that the UUV maintains enough thrust to support the timely completion of the mission. The thrust monitor may change the system configuration dynamically (i.e., turning on additional thrusters when the actual thrust falls below the expected thrust.)
Figure 3. A screenshot of the Gazebo UUV simulator.
Environmental Anomalies. Next, we randomly injected abnormal environmental events throughout the operation of the UUV. We list two failures that we investigated:
@@ 6.1. Simulator
  • Loss of visual contact with the underwater pipeline: The low water visibility makes it difficult for the UUV to detect and follow the underwater pipeline. When it occurs during the pipeline inspection, it may result in the UUV losing sight of the pipeline and requiring it to dive deeper in the short term, regain the visual line of sight, and continue the inspection progress. The change in the water visibility is introduced randomly during the simulation.
@@
  • Thruster failure: A thruster failure causes the engine of the UUV to provide partial propulsion, move erratically, or shut down completely. Similarly, thruster failures are modeled and injected in a stochastic manner.
6.2. Environment Model
@@
Our framework leverages a model of the environment during the adaptation process. For the UUV system, the environmental model captures the (1) physical dynamics of the system (2) thruster estimation from the engine, and (3) the estimated coordinates of the seabed pipeline. The dynamics model is used to estimate the velocity of the UUV based on the accelerating or decelerating actions. The thruster estimation is based on the configuration of the thrusters (i.e., which ones are turned on or off). Lastly, the estimated coordinates of the seabed pipeline are based on information received via onboard sensors. The coordinate information is also used to keep track of the inspection progress and guide the diving operation toward the pipeline.
@@
All aspects of the environmental model are specified in the MiniZinc modeling language, translated to linear constraints and solved using MILP during the adaptation process.
7. Evaluation
@@
This section describes the evaluation of our proposed approach. We present the following:
@@
  • RQ1: Does our approach achieve a higher overall system utility than existing state-of-the-art approaches?
@@
7.1. Experimental Design
We conducted a set of experiments to evaluate the proposed adaptation approach through comparison against with state-of-the-art self-adaptation method in TOMASys (Bermejo-Alonso et al., 2016).To ensure that our adaptation approach generalizes across various scenarios, we randomly generated 100 different system setups and failure scenarios, including the duration of the simulation, times when failures are injected, changes in the water visibility, number of usable thrusters, the initial position of the UUV, and the location of the underwater pipeline.
@@ 7.1. Experimental Design
To compare our proposed method and the baseline approach, we use the cumulative utility as the metric, which is measured by calculating the robustness of the minimal requirement for the signal collected during the degradation and recovery process. The reason why we chose to measure the satisfaction of the minimal requirement is that it is the lower bound of the instantiated STL formula under which the system feature ceases to function and provides useful utility. We assume that the robustness metric is a suitable medium to reflect the desirability of the system behavior. The robustness is measured upon the start of the degradation events and ends upon the satisfaction of the optimal requirement ($\varphi_{opt}$) which indicates the end of the recovery. If the optimal requirement is never reached due to persistent environmental disruptions, the cumulative robustness measurement will continue until the end of the simulation life cycle.
@@
Before we conducted the experiments for the case studies, we created the following hypotheses to be tested:
@@
  • H1: Our approach results in a higher cumulative utility throughout degradation and recovery processes than the existing method.
@@
  • H2: Our approach results in a higher run-time overhead but it does not disrupt normal system operations.
@@
All our experiments were run on a Ubuntu desktop machine with 16 GB RAM, 6-core Intel Core i5, and a NVIDIA GeForce RTX 3060 graphics card.
7.2. Seabed Pipe Inspection Case Study
Recall the example mentioned in Section 2, where a UUV is conducting a mission to inspect underwater pipes on the seabed. The optimal STL formula $\varphi_{opt}$ and the minimal requirement $\varphi_{min}$ for the water visibility and thrust requirement are specified as follows:
@@ 7.2. Seabed Pipe Inspection Case Study
(20) $\begin{split}&\varphi_{visibility\_opt}:\Box_{[0,1]}(\texttt{visibility}<20\Rightarrow\\ &\quad\quad\quad\quad\quad\Diamond_{[0,5]}\texttt{distance\_to\_pipeline}<10)\end{split}$
@@
(21) $\begin{split}&\varphi_{visibility\_min}:\Box_{[0,1]}(\texttt{visibility}<20\Rightarrow\\ &\quad\quad\quad\quad\quad\Diamond_{[0,15]}\texttt{distance\_to\_pipeline}<10)\end{split}$
(22) $\begin{split}&\varphi_{thruster\_opt}:\Box_{[0,1]}(\texttt{thrust}>100)\end{split}$
(23) $\begin{split}&\varphi_{thruster\_min}:\Box_{[0,1]}(\texttt{thrust}>50)\end{split}$
The visibility requirement $\varphi_{visibility}$ ensures that the UUV can closely observe the pipelines upon the visibility falling below the safety threshold 20. The thrust requirements allow the system to continuously provide adequate propulsion to ensure a timely mission completion.
@@
Figure 4. Cumulative robustness for visibility and thruster monitor features for the pipeline inspection case study
7.3. Experiment Results
@@
In Fig 4, charts (A) and (B) show the cumulative robustness values for both the visibility and thruster monitor feature while the system is encountering thruster failures or low water visibility events. The results show that our adaptation approach achieves a higher cumulative robustness value than the baseline self-adaptive approach. Specifically, in the case of the water visibility monitor, it achieves an approximately 2-fold increase in cumulative robustness (with a 26.9 increase in robustness). In the case of the thruster monitor feature, our approach has a 24.3 lead in cumulative robustness.
@@
The SUAVE artifact (Silva et al., 2023) also comes with a set of mission-related metrics that measure the quality of the mission completion. We reused some of the metrics, namely, the distance of the pipeline inspected. We discovered that our approach, on average, can inspect 86.10 meters of the pipeline while the baseline approach can only inspect 49.63 meters, an increase of about 74%. However, the standard deviation of our approach (36.68) is slightly worse than the baseline approach (23.30), meaning that the actual performance fluctuates more across various failure scenarios in the 100 experiment runs.
@@
There are two main reasons for this increase in the cumulative utility and the quality of the mission completion. First, it is the dynamic nature of the action planning based on the specific scenario or context. For instance, TOMASys relies on a fixed set of actions (determined at design time) to address each scenario. To satisfy the objective of regaining contact with the pipeline after losing visuals, TOMASys simply has a predefined action to dive several meters down the seabed, whereas our approach predicts all possible signals given available actions using an environmental model. Then it chooses the actions that result in the most desirable signals. Second, it is the ability to flexibly adapt goals in different situations—namely, adjusting the requirements when they are not attainable or have the potential for improvements. The dynamic action planning and the requirement adjustments result in a more measurable system utility, and therefore, more robust and desirable system actions.
7.4. Performance Overhead
@@
Since the requirement adaptation approach requires the use of MILP to search for requirements and plan actions at run time, it has a higher overhead than the baseline approach that uses actions that are defined at design time. We measure the overhead as the average additional time the system uses per control cycle as a result of deploying our adaptation approach compared to the baseline approach using TOMASys. Consequently, we measured the average overhead across both the visibility and thrust monitor features as 0.35 seconds. We have not observed noticeable delays or disruptions to our UUV during the operation, as the UUV controller was running at 2Hz, resulting in a window of 0.5 seconds for each cycle of the controller update.
@@
7.5. Threats to Validity
Internal Validity: The sampled configuration space is partial and may not capture all exceptional scenarios exhaustively. However, we have mitigated selection biases by randomizing the configuration parameters. For example, at which point failures occur, the initial location of the UUV, the setup of the underwater pipelines, etc. In addition, we are using a deterministic environmental model that captures simple physical dynamics. This is to simplify the scope of the problem and reduce the search space for the changed requirements and corresponding actions. However, the model may not capture the dynamic of the real world caused by external factors (i.e., UUVs may deviate from the direction it is accelerating towards due to water currents). We attempted to mitigate this discrepancy by manually inspecting the source code of the ArduSub simulator to ensure the model largely captures the logic of the simulator.
@@ 7.5. Threats to Validity
External Validity: The overhead of the simulation is application-specific and hardware-dependent. More restrictive hardware or outdated MILP solver may increase the runtime overhead of our framework and therefore require more performance tuning than our existing software artifacts.
Furthermore, the result of the case study may not be generalized to all applications of CPS. Depending on the specific applications (i.e., electricity grids, UAVs), the system requirement may be expressed differently, and so does the way to degrade the systems. However, we believe the concept of using requirement weakening and strengthening to perform degradation and recovery is requirement-agnostic, and that it can be generalized across different domains.
8. Related Work
@@
Graceful Degradation and Recovery. Graceful degradation is the ability of a system to maintain an acceptable level of functionality even when a significant portion of the system is rendered inoperable due to environmental disturbances. Recovery refers to the ability of the system to restore to a safe or desired state after a failure. Both areas are well-studied in control systems and human-machine interfaces. Many degradation frameworks have been developed for different applications and domains, such as adaptive cruise control (Ploeg et al., 2015), autonomous drones (Chu et al., 2023; Zudaire et al., 2022), software user interfaces (Florins and Vanderdonckt, 2004), industrial control systems (Buckworth et al., 2023; D’Ippolito et al., 2014), and design patterns for degradation(Saridakis, 2009), respectively. There is also a large body of work on system recovery. (Chen et al., 2022) proposed an STL-based resilient framework that enables analysis of trade-offs between time-to-recovery and durability in a cruise control system using multi-objective optimization, while (Kong et al., 2018) and (Venkitakrishnan, 2002) propose checkpointing techniques (i.e., storing system traces with a finite horizon leading up to the current state). Lastly, (Jagtap et al., 2020) and (Abad et al., 2016) use system restart and reset to recover the system from a failure. As far as we know, no prior work provides a coordination mechanism to facilitate these two processes in a single system.
@@
Self-adaptive Systems. Self-adaptive systems aim to design systems that are capable of adjusting to changes in the environment. The most influential reference control model in autonomic and self-adaptive systems is the MAPE-K (Arcaini et al., 2015) architecture, which stands for Monitor-Analyze-Plan-Execute over a shared Knowledge, on which our resolution architecture is based. Self-healing systems (Shaw, 2002) have been proposed by Shaw, et al. to address the problem of coping with fluctuations and uncertainties in the environment. The idea is that the system always keeps a background maintenance process running regardless of the current state of the system and the environment (i.e. garbage collection, optimized network routing, etc.). This way, the system will eventually maintain internal stability and continuous improvement despite external changes. This state is also known as a homeostasis state. Our approach draws inspiration from this idea, in that we attempt to ensure that the system aims to achieve the most desirable requirement relative to the current requirement.
The concept of meta-control has been used for design-time analysis. For example, OpenODD (Reich et al., 2023) has been used as a scenario-based analysis framework for autonomous systems, especially in self-driving cars. TOMASys (Bermejo-Alonso et al., 2016) provides a comprehensive set of reconfiguration strategies ranging from architectural reconfiguration, to monitor construction, and then to scenario-based definition of adaptation tactics.
Requirement Adaptation. Requirement adaptation has been investigated in the context of self-adaptive systems. The idea of weakening a requirement has been studied for feature interaction resolution (Chu et al., 2023), handling a violation of environmental assumptions and safety properties (Buckworth et al., 2023; Zudaire et al., 2022), controller synthesis (Cavezza et al., 2018; Buyukkocak and Aksaray, 2022), and goal adjustment for security-critical systems (Tun et al., 2021). As far as we know, however, these existing works focus on how to gracefully degrade the system using requirement relaxation, instead of combining both degradation and recovery stages or attempting to coordinate them using requirement adaptation.
RELAX (Whittle et al., 2010) is designed to support expressing requirements that explicitly capture uncertainties about runtime system behavior based on fuzzy temporal logic. For example, with RELAX, users can specify requirements like, ”Once a user request is sent, it should be processed as early as possible”. Although our approach differs in that it relies on STL as the underlying specification formalism, an interesting future work would be to investigate the utility of RELAX for the type of degradation-recovery loop that we have explored.
@@
Both (Wohlrab et al., 2022) and (Li et al., 2023) study goal adaptation in the context of changing stakeholder goals or user preferences. However, their proposed methods are designed to be dependent on human inputs, not for automatically adapting to changing environments. (Bennaceur et al., 2019) focuses on requirement adaptation based on resource constraints and proposes a 2-tier adaptation framework that (1) first tries to fulfill goals by finding alternative resources and (2) if needed, deviates the goal slightly from the original one to compensate for limited resources. Our approach differs from these approaches in that our utility is represented as the desirability of requirements over run-time signals, instead of adequate resources being allocated.
9. Limitations and Future Work